# Near-extremal hydrodynamics and the holographic product formula

### v1

@@ 1 Introduction and summary
Most finite temperature field theories are described by hydrodynamics over long times and distances. That is, for energies $\omega$ and momenta $k$ much smaller than temperature $T$, the dynamics of most systems reduces to the conservation equations for their conserved currents. This includes in particular the majority of holographic theories, whose gravitational dual should therefore also admit a fluid description in the hydrodynamic regime. This result was established explicitly in Bhattacharyya:2007vjd, and has been called the fluid/gravity correspondence.
A more surprising result that has since been observed in holographic systems at finite density, is that hydrodynamics seemingly continues to hold in the low temperature limit $T\ll\mu$ (with $\mu$ the chemical potential), as long as energies and momenta remain much smaller than $\mu$ Edalati:2010hk; Edalati:2010pn; Moitra:2020dal; Davison:2013bxa; Davison:2022vqh; Jarvinen:2023xrx; Arnaudo:2024sen. In particular, at leading order in the low $\omega,k$ expansion, holographic two-point functions were found to agree with the zero-temperature limit of hydrodynamic correlators Davison:2013bxa; Moitra:2020dal. Not only do the hydrodynamic poles survive, but no sign was found at this order of the branch cut that is known to arise at $T=0$ Faulkner:2009wj; DHoker:2010xwl, which is related to the lower-dimensional conformal field theory (CFT) emerging in the infra-red (IR) of these holographic systems.
The recent calculation of Gouteraux:2025kta at next-to-leading order made it clear that the branch cut puzzle only exists at leading order, since the next order already contains logarithmic signatures of the branch cut. However, the result of Gouteraux:2025kta still does not contain a clear imprint from the IR CFT.
@@
In this work, we exploit the special analytic properties of holographic correlators to get more insight into their structure in the near-extremal hydrodynamic regime $\omega,k,T\ll\mu$. Specifically, we derive the consequences of the recently established holographic product formula Dodelson:2023vrw, which makes it possible to express holographic spectral functions as a product over their poles. This formula therefore allows to write the general form of near-extremal holographic correlators based on their known pole structure, which features three types of poles (see figure 1): the hard poles with gaps of order $\mathcal{O}(\mu)$, the soft poles with gaps of order $\mathcal{O}(T)$ (related to the IR CFT), and the gapless poles, which behave like hydrodynamic modes at leading order in the near-extremal hydrodynamic limit. From this structure, we deduce the near-extremal hydrodynamic product formula
@@
with $\rho$ the spectral function, $\mathcal{G}_{s}$ the product over the soft poles and $\gamma$ a generalized susceptibility, which is and analytic in $\omega$ for $\omega\ll\mu$.
@@
In some cases, the soft poles can be shown (numerically) to precisely approach the IR poles in the near-extremal limit, up to corrections of order $\mathcal{O}(T/\mu)$ Edalati:2010hk; Davison:2022vqh. In these cases, the near-extremal product formula simplifies at leading order in $T/\mu$ to
@@
with $\mathcal{G}_{\text{IR}}$ the thermal correlator in the IR CFT, and $\tilde{\gamma}$ analytic in $\omega$. This expression is valid in the general near-extremal hydrodynamic regime, including when $\omega$ is of the same order as the temperature.
@@
The rest of this paper is organized as follows. Section 2 discusses in more details the developments summarized above, whereas section 3 presents several examples where the product formula can be used to determine the general structure of near-extremal spectral functions. This includes probe gauge field (section 3.1) and helicity-0111The helicity-1 shear-channel of metric perturbations has been analyzed in the recent work Bhattacharya:2025vyi. metric (section 3.2) perturbations on the AdS5-Reissner-Nordström background – with a (0+1)-dimensional CFT in the IR – as well as scalar perturbations of magnetized black branes (section 3.3), whose IR description is in terms of a (1+1)-dimensional CFT DHoker:2009mmn; DHoker:2010xwl. As intermediate results, we computed numerically the low-lying poles of the spectral functions in the near-extremal regime for all three cases. These numerical results are presented in appendices D, G and H.
Section 4 discusses a concrete application of the product formula to the problem analyzed in Jarvinen:2023xrx, related to neutrino transport in dense holographic matter. Improving the hydrodynamic approximation to the weak-current correlators with the IR scaling is shown to give a better description of the neutrino opacity near the Fermi surface (see figure 3).
Several additional details are provided in the appendices. In particular, in appendix B, we show how the next-to-leading order result of Gouteraux:2025kta is consistent with the product formula, which completes it at low energy $\omega$ by having the IR power-law appear explicitly.
@@ 2 The product formula
We present in this section the product formula for holographic two-point functions, as introduced in Dodelson:2023vrw. We first review the formula and its conditions of application, and then work out its consequences for the structure of holographic correlators in the near-extremal regime.
@@ 2.1 The formula
The product formula Dodelson:2023vrw is a statement about thermal two-point functions at equilibrium. For a given operator $\mathcal{O}$, it is best expressed in terms of the two-sided correlator $G_{12}$, defined in position space as
@@
implying that $1/G_{12}(\omega)$ is an entire function of order 1 Dodelson:2023vrw.
Meromorphicity and the asymptotic behavior (5) are generic properties of holographic correlators Hartnoll:2005ju; Festuccia:2005pi; Festuccia:thesis. The property that is not generically obeyed is for the correlator to be free of zeroes. It can be shown to be the case upon two conditions Dodelson:2023vrw:
@@
    A given set of decoupled fluctuation equations can always be written in the Schrödinger form (6) upon appropriate field redefinitions (see e.g. appendix B of Arean:2013tja). However, finding variables that obey decoupled equations is more difficult in general, and not always possible [^5] Another possible issue may be that, even though the product formula applies to the correlators dual to the decoupled variables, it does not carry over to the original correlators of interest. This typically arises when the relation between the two types of correlators is non-linear, as in the example presented in appendix F.. A non-trivial example of such variables is given by the “master fields” for the metric and gauge fields fluctuations of the RN background Kodama:2003kk; Edalati:2010pn; Edalati:2010hk;
  • The potential $V(z)$ should be regular for all $z>0$, with boundary ($z\to 0$) and horizon ($z\to\infty$) asymptotics obeying
    $V(z)\underset{z\to 0}{\sim}\frac{\nu^{2}-\frac{1}{4}}{z^{2}}\quad,\quad z^{2}V(z)\underset{z\to\infty}{\to}0\,$, (8)
    with $\nu^{2}\geq 0$. In practice, all regular holographic solutions obey the IR condition at finite temperature, with $V(z)\sim\mathrm{e}^{-4\pi z/\beta}$ for $z\to\infty$ Dodelson:2023vrw. The UV $z^{-2}$ behavior of the potential is also generic, although in some cases the parameter $\nu^{2}$ may become negative. So the effective condition on the potential is the absence of singularities in the bulk – i.e. for all $z\in(0,\infty)$ – together with $\nu^{2}\geq 0$.
Note that, even though it will be important for the developments of this work that the fluctuation equations decouple, it is most of the time not crucial that $G_{12}$ is actually free of zeroes. In other words, most of our results still apply for a Schrödinger potential that admits singularities. In this case the generalized susceptibilities introduced below may admit zeroes. There is one exception in subsection 2.3 where we mention explicitly that the no-zeroes conditions should hold.
@@ 2.2 Application to near-extremal hydrodynamics
We now consider a thermal state with a dimensionful source parameter $\mu$, in the near-extremal limit $T\ll\mu$, such that the bulk dual corresponds to a near-extremal black brane with AdSp-Schwarzschild near-horizon geometry. The low energy excitations of such states are described by a CFTp-1 at each point in the transverse space of dimension dp+1d-p+1 Faulkner:2009wj; DHoker:2009mmn. In momentum space, this translates into a family of (p1)(p-1)-dimensional CFT’s labeled by the transverse momentum $\vec{k}_{d-p+1}$. Examples with $p=2$ include the charged AdS Reissner-Nordström (RN) black brane, with $\mu$ the chemical potential, and the neutral solution of Bardoux:2012aw; Andrade:2013gsa, with $\mu$ the source for the translation-symmetry breaking axions. The magnetized CFT2 of DHoker:2009mmn is an example with $p=3$, where $\mu$ is related to the magnetic field via $\mu=\sqrt{B}$.
@@
  • Gapless poles $\omega^{g}_{n}(\vec{k})$, which go to zero at zero momentum. Such poles were exhibited in Edalati:2010hk; Edalati:2010pn; Davison:2013bxa; Moitra:2020dal. They do not go to zero in the limit of vanishing temperature [^6] With this definition, the results of Davison:2022vqh indicate that the gapless poles (and the soft poles) are generically discontinuous as a function of momentum near a collision with a soft pole. The discontinuities are small, of order $\mathcal{O}(r_{e}T)$., where they instead follow a dispersion relation which at leading order in momentum is similar to hydrodynamic modes [^7] Because of this similarity, and following the terminology of Blake:2017ris, a possible name for these modes would be “quantum hydrodynamic modes”. They clearly differ from standard hydrodynamic modes since they contain non-analytic logarithmic terms at higher order in $r_{e}|\vec{k}|$ Gouteraux:2025kta. (e.g. a sound mode $\omega^{g}(\vec{k})\sim c_{s}|\vec{k}|-i\Gamma\vec{k}^{2}+\mathcal{O}(r_{e}^{3}\vec{k}^{4})$ Edalati:2010pn; Moitra:2020dal or diffusive mode $\omega^{g}(\vec{k})\sim-iD\vec{k}^{2}+\mathcal{O}(r_{e}^{3}\vec{k}^{4})$ Edalati:2010hk; Davison:2013bxa; Moitra:2020dal).
@@
where $\tilde{\gamma}_{e}$ is analytic in $\omega$ for $|\omega|,|\vec{k}|\ll r_{e}^{-1}$, and we defined the soft factor as
@@
At this point it is interesting to consider the limit $T\to 0$, for which the structure of the soft sector (15) is expected to simplify. For $|\omega|,|\vec{k}|\ll r_{e}^{-1}$ and away from gapless poles, holographic near-extremal correlators computed from (6) can indeed be shown to be proportional to the corresponding IR correlators Faulkner:2009wj; DHoker:2010xwl [^8] We generalize the results of Faulkner:2009wj; DHoker:2010xwl to general near-extremal black holes in appendix A. Here we are considering the regime where not only $r_{e}|\omega|$ but also $r_{e}|\vec{k}|$ is small.
@@
where the $\sinh$ comes from (3), and we split the spatial momentum into the part that is along the CFT direction $\vec{k}_{p-2}$ and the transverse part $\vec{k}_{d-p+1}$. $\Delta(\vec{k}_{d-p+1})$ is the IR conformal dimension associated to $G_{12}$. For generic values of the momentum, $\Delta$ is not an integer, which implies that the correlator $G_{12}$ has branch points at $\omega=\pm|\vec{k}_{p-2}|$ (there is a single branch point at $\omega=0$ for $p=2$). This means that the soft poles merge into branch cuts in the zero-temperature limit Faulkner:2009wj.
@@
$\text{Im}G_{R}(\omega,\vec{k})=\frac{\gamma_{e}(\omega,\vec{k})\big(r_{e}^{2}(\omega^{2}-\vec{k}_{p-2}^{2})\big)^{\Delta(\vec{k}_{d-p+1})-(p-1)/2}}{\prod_{n\in\text{hydro}}\left(\omega^{2}-\omega_{n}^{h}(\vec{k})^{2}\right)\left(\omega^{2}-(\omega_{n}^{h*}(\vec{k}))^{2}\right)}\,.$ (19)
@@
In particular, it is an important aspect of (19) that it allows to extend the near-extremal hydrodynamic expansion to the region near the branch points. In a standard expansion, the power-law (18) would indeed be expanded at low momentum, generating terms of the form [^9] An example is given by the result of Gouteraux:2025kta discussed in appendix B, which contains the first term in this expansion. $\vec{k}_{d-p+1}^{2m}\log{(\omega^{2}-\vec{k}_{p-2}^{2})}$, which imply a breakdown of the expansion at the branch points. The leading behavior of the numerator (18) resums all such terms, which allows to include the branch points in the near-extremal hydrodynamic regime. Section 4 gives a concrete example of this IR improvement in the case of $p=2$.
Another comment is related to the number of gapless poles that appear in the denominator of (14), which may actually depend on the order of the hydrodynamic expansion. Appendix B gives an example of this, where the single leading order diffusive mode $\omega_{D}(\vec{k})=-iD\vec{k}^{2}$, splits into two poles with finite real parts at next-to-leading order Gouteraux:2025kta. In this case $\gamma_{e}(\omega,\vec{k})$ should have zeroes at the appropriate order to account for the pole merging.

### v2

@@ 1 Introduction and summary
Most finite temperature field theories are described by hydrodynamics over long times and distances. That is, for energies $\omega$ and momenta $k$ much smaller than temperature $T$, the dynamics of most systems reduces to the conservation equations for their conserved currents. This includes in particular the majority of holographic theories, whose gravitational dual should therefore also admit a fluid description in the hydrodynamic regime. This result was established explicitly in Bhattacharyya et al. (2008), and has been called the fluid/gravity correspondence.
A more surprising result that has since been observed in holographic systems at finite density, is that hydrodynamics seemingly continues to hold in the low temperature limit $T\ll\mu$ (with $\mu$ the chemical potential), as long as energies and momenta remain much smaller than $\mu$ Edalati et al. (2010b, a); Moitra et al. (2021); Davison and Parnachev (2013); Davison et al. (2023); Järvinen et al. (2023); Arnaudo and Withers (2025). In particular, at leading order in the low $\omega,k$ expansion, holographic two-point functions were found to agree with the zero-temperature limit of hydrodynamic correlators Davison and Parnachev (2013); Moitra et al. (2021). Not only do the hydrodynamic poles survive, but no sign was found at this order of the branch cut that is known to arise at $T=0$ Faulkner et al. (2011); D’Hoker et al. (2011), which is related to the lower-dimensional conformal field theory (CFT) emerging in the infra-red (IR) of these holographic systems.
The recent calculation of Goutéraux et al. (2025) at next-to-leading order made it clear that the branch cut puzzle only exists at leading order, since the next order already contains logarithmic signatures of the branch cut. However, the result of Goutéraux et al. (2025) still does not contain a clear imprint from the IR CFT.
@@
In this work, we exploit the special analytic properties of holographic correlators to get more insight into their structure in the near-extremal hydrodynamic regime $\omega,k,T\ll\mu$. Specifically, we derive the consequences of the recently established holographic product formula Dodelson et al. (2024), which makes it possible to express holographic spectral functions as a product over their poles. This formula therefore allows to write the general form of near-extremal holographic correlators based on their known pole structure, which features three types of poles (see figure 1): the hard poles with gaps of order $\mathcal{O}(\mu)$, the soft poles with gaps of order $\mathcal{O}(T)$ (related to the IR CFT), and the gapless poles, which behave like hydrodynamic modes at leading order in the near-extremal hydrodynamic limit. From this structure, we deduce the near-extremal hydrodynamic product formula
@@
with $\rho$ the spectral function, $\mathcal{G}_{s}$ the product over the soft poles and $\gamma$ a generalized susceptibility, which is analytic for $\omega\ll\mu$.
@@
In some cases, the soft poles can be shown (numerically) to precisely approach the IR poles in the near-extremal limit, up to corrections of order $\mathcal{O}(T/\mu)$ Edalati et al. (2010b); Davison et al. (2023). In these cases, the near-extremal product formula simplifies at leading order in $T/\mu$ to
@@
with $\mathcal{G}_{\text{IR}}$ the thermal correlator in the IR CFT, and $\tilde{\gamma}$ analytic. This expression is valid in the general near-extremal hydrodynamic regime, including when $\omega$ is of the same order as the temperature.
@@
The rest of this paper is organized as follows. Section 2 discusses in more details the developments summarized above, whereas section 3 presents several examples where the product formula can be used to determine the general structure of near-extremal spectral functions. This includes probe gauge field (section 3.1) and helicity-0111The helicity-1 shear-channel of metric perturbations has been analyzed in the recent work Bhattacharya et al. (2025). metric (section 3.2) perturbations on the AdS5-Reissner-Nordström background – with a (0+1)-dimensional CFT in the IR – as well as scalar perturbations of magnetized black branes (section 3.3), whose IR description is in terms of a (1+1)-dimensional CFT D’Hoker and Kraus (2009); D’Hoker et al. (2011). As intermediate results, we computed numerically the low-lying poles of the spectral functions in the near-extremal regime for all three cases. These numerical results are presented in appendices D, G and H.
Section 4 discusses a concrete application of the product formula to the problem analyzed in Järvinen et al. (2023), related to neutrino transport in dense holographic matter. Improving the hydrodynamic approximation to the weak-current correlators with the IR scaling is shown to give a better description of the neutrino opacity near the Fermi surface (see figure 3).
Several additional details are provided in the appendices. In particular, in appendix B, we show how the next-to-leading order result of Goutéraux et al. (2025) is consistent with the product formula, which completes it at low energy $\omega$ by having the IR power-law appear explicitly.
@@ 2 The product formula
We present in this section the product formula for holographic two-point functions, as introduced in Dodelson et al. (2024). We first review the formula and its conditions of application, and then work out its consequences for the structure of holographic correlators in the near-extremal regime.
@@ 2.1 The formula
The product formula Dodelson et al. (2024) is a statement about thermal two-point functions at equilibrium. For a given operator $\mathcal{O}$, it is best expressed in terms of the two-sided correlator $G_{12}$, defined in position space as
@@
implying that $1/G_{12}(\omega)$ is an entire function of order 1 Dodelson et al. (2024).
Meromorphicity and the asymptotic behavior (5) are generic properties of holographic correlators Hartnoll and Kumar (2005); Festuccia and Liu (2006); Festuccia (2009). The property that is not generically obeyed is for the correlator to be free of zeroes. It can be shown to be the case upon two conditions Dodelson et al. (2024):
@@
    A given set of decoupled fluctuation equations can always be written in the Schrödinger form (6) upon appropriate field redefinitions (see e.g. appendix B of Areán et al. (2013)). However, finding variables that obey decoupled equations is more difficult in general, and not always possible [^5] Another possible issue may be that, even though the product formula applies to the correlators dual to the decoupled variables, it does not carry over to the original correlators of interest. This typically arises when the relation between the two types of correlators is non-linear, as in the example presented in appendix F.. A non-trivial example of such variables is given by the “master fields” for the metric and gauge fields fluctuations of the RN background Kodama and Ishibashi (2004); Edalati et al. (2010a, b);
  • The potential $V(z)$ should be regular for all $z>0$, with boundary ($z\to 0$) and horizon ($z\to\infty$) asymptotics of the form
    $V(z)\underset{z\to 0}{\sim}\frac{\nu^{2}-\frac{1}{4}}{z^{2}}\quad,\quad V(z)=\sum_{n\geq 1}\mathrm{e}^{-\frac{4\pi n}{\beta}z}\,$, (8)
    with $\nu^{2}\geq 0$. In practice, all regular holographic solutions obey the IR condition at finite temperature Dodelson et al. (2024). The UV $z^{-2}$ behavior of the potential is also generic, although in some cases the parameter $\nu^{2}$ may become negative. So the effective condition on the potential is the absence of singularities in the bulk – i.e. for all $z\in(0,\infty)$ – together with $\nu^{2}\geq 0$.
Note that, even though it will be important for the developments of this work that the fluctuation equations decouple, it is most of the time not crucial that $G_{12}$ is actually free of zeroes. In other words, most of our results are expected to still apply for a Schrödinger potential that admits singularities. In this case the generalized susceptibilities introduced below may admit zeroes. There is one exception in subsection 2.3 where we mention explicitly that the no-zeroes conditions should hold.
@@ 2.2 Application to near-extremal hydrodynamics
We now consider a thermal state with a dimensionful source parameter $\mu$, in the near-extremal limit $T\ll\mu$, such that the bulk dual corresponds to a near-extremal black brane with AdSp-Schwarzschild near-horizon geometry. The low energy excitations of such states are described by a CFTp-1 at each point in the transverse space of dimension dp+1d-p+1 Faulkner et al. (2011); D’Hoker and Kraus (2009). In momentum space, this translates into a family of (p1)(p-1)-dimensional CFT’s labeled by the transverse momentum $\vec{k}_{d-p+1}$. Examples with $p=2$ include the charged AdS Reissner-Nordström (RN) black brane, with $\mu$ the chemical potential, and the neutral solution of Bardoux et al. (2012); Andrade and Withers (2014), with $\mu$ the source for the translation-symmetry breaking axions. The magnetized CFT2 of D’Hoker and Kraus (2009) is an example with $p=3$, where $\mu$ is related to the magnetic field via $\mu=\sqrt{B}$.
@@
  • Gapless poles $\omega^{g}_{n}(\vec{k})$, which go to zero at zero momentum. Such poles were exhibited in Edalati et al. (2010b, a); Davison and Parnachev (2013); Moitra et al. (2021). They do not go to zero in the limit of vanishing temperature [^6] With this definition, the results of Davison et al. (2023) indicate that the gapless poles (and the soft poles) are generically discontinuous as a function of momentum near a collision with a soft pole. The discontinuities are small, of order $\mathcal{O}(r_{e}T)$., where they instead follow a dispersion relation which at leading order in momentum is similar to hydrodynamic modes [^7] Because of this similarity, and following the terminology of Blake et al. (2018), a possible name for these modes would be “quantum hydrodynamic modes”. They clearly differ from standard hydrodynamic modes since they contain non-analytic logarithmic terms at higher order in $r_{e}|\vec{k}|$ Goutéraux et al. (2025). (e.g. a sound mode $\omega^{g}(\vec{k})\sim c_{s}|\vec{k}|-i\Gamma\vec{k}^{2}+\mathcal{O}(r_{e}^{3}\vec{k}^{4})$ Edalati et al. (2010a); Moitra et al. (2021) or diffusive mode $\omega^{g}(\vec{k})\sim-iD\vec{k}^{2}+\mathcal{O}(r_{e}^{3}\vec{k}^{4})$ Edalati et al. (2010b); Davison and Parnachev (2013); Moitra et al. (2021)).
@@
where $\tilde{\gamma}_{e}$ is analytic for $|\omega|,|\vec{k}|\ll r_{e}^{-1}$, and we defined the soft factor as
@@
At this point it is interesting to consider the limit $T\to 0$, for which the structure of the soft sector (15) is expected to simplify. For $|\omega|,|\vec{k}|\ll r_{e}^{-1}$ and away from gapless poles, holographic near-extremal correlators computed from (6) can indeed be shown to be proportional to the corresponding IR correlators Faulkner et al. (2011); D’Hoker et al. (2011) [^8] We generalize the results of Faulkner et al. (2011); D’Hoker et al. (2011) to general near-extremal black holes in appendix A. Here we are considering the regime where not only $r_{e}|\omega|$ but also $r_{e}|\vec{k}|$ is small.
@@
where the $\sinh$ comes from (3), and we split the spatial momentum into the part that is along the CFT direction $\vec{k}_{p-2}$ and the transverse part $\vec{k}_{d-p+1}$. $\Delta(\vec{k}_{d-p+1})$ is the IR conformal dimension associated to $G_{12}$. For generic values of the momentum, $\Delta$ is not an integer, which implies that the correlator $G_{12}$ has branch points at $\omega=\pm|\vec{k}_{p-2}|$ (there is a single branch point at $\omega=0$ for $p=2$). This means that the soft poles merge into branch cuts in the zero-temperature limit Faulkner et al. (2011).
@@
$\text{Im}G_{R}(\omega,\vec{k})=\frac{\gamma_{e}(\omega,\vec{k})\big(r_{e}^{2}(\omega^{2}-\vec{k}_{p-2}^{2})\big)^{\Delta(\vec{k}_{d-p+1})-(p-1)/2}}{\prod_{n\in\text{gapless}}\left(\omega^{2}-\omega_{n}^{g}(\vec{k})^{2}\right)\left(\omega^{2}-(\omega_{n}^{g*}(\vec{k}))^{2}\right)}\,.$ (19)
@@
In particular, it is an important aspect of (19) that it allows to extend the near-extremal hydrodynamic expansion to the region near the branch points. In a standard expansion, the power-law (18) would indeed be expanded at low momentum, generating terms of the form [^9] An example is given by the result of Goutéraux et al. (2025) discussed in appendix B, which contains the first term in this expansion. $\vec{k}_{d-p+1}^{2m}\log{(\omega^{2}-\vec{k}_{p-2}^{2})}$, which imply a breakdown of the expansion at the branch points. The leading behavior of the numerator (18) resums all such terms, which allows to include the branch points in the near-extremal hydrodynamic regime. Section 4 gives a concrete example of this IR improvement in the case of $p=2$.
Another comment is related to the number of gapless poles that appear in the denominator of (14), which may actually depend on the order of the hydrodynamic expansion. Appendix B gives an example of this, where the single leading order diffusive mode $\omega_{D}(\vec{k})=-iD\vec{k}^{2}$, splits into two poles with finite real parts at next-to-leading order Goutéraux et al. (2025). In this case $\gamma_{e}(\omega,\vec{k})$ should have zeroes at the appropriate order to account for the pole merging.